OpenStudy (anonymous):

rationalize the denominator: 9 / (7-3sqrt3) - (6-4sqrt3)

OpenStudy (anonymous):

ive narrowed the bottom down to either -2, 3+sqrt3, or 4

hero (hero):

\(\large\frac{9}{(7-3\sqrt{3}) - (6 - 4\sqrt{3})}\)

OpenStudy (anonymous):

yes thats the problem

hero (hero):

\(\frac{9}{1+\sqrt{3}}\)

hero (hero):

Multiply top and bottom by the expression in the denominator

OpenStudy (anonymous):

the top would be 9+9sqrt3?

hero (hero):

Yeah

hero (hero):

wait

hero (hero):

I think you have to multiply top and bottom by the conjugate

hero (hero):

that's the only way to get rid of the radical

hero (hero):

So muliply top and bottom by \(1 - \sqrt{3}\)

hero (hero):

That way, you'll end up with: \(\frac{9 - 9\sqrt{3}}{1-3}\)

hero (hero):

denominator would simplify to -2